The market interest rate is 9 percent and is expected to stay at that level. Consumers can borrow and lend all they want at this rate. Would you prefer a ​$150 gift today or a ​$155 gift next​ year? You would ▼______ .(prefer the gift today/ prefer the gift next year /be indifferent .)

What if the interest rate were 15 ​percent? If the interest rate were 15 ​percent, then you would ________▼ (prefer the gift today/ prefer the gift next year /be indifferent .)
Answer #1

Present value of $150 gift today = $150

Present value of $155 gift next year = 155/1.09 = 142.20

So, you would prefer the gift today.

If interest rate were 15 percent,

PV of $150 gift today = $150

PV of $155 gift next year = 155/1.15 = 134.78

So, you would prefer the gift today.